# VITA: Zero-Shot Value Functions via Test-Time Adaptation of Vision–Language Models

### v4

@@ Abstract
Vision–Language Models (VLMs) show promise as zero-shot goal-conditioned value functions, but their frozen pre-trained representations limit generalization and temporal reasoning. We introduce VITA, a zero-shot value function learning method that enhances both capabilities via test-time adaptation. At inference, a lightweight adaptation module is updated via a gradient step on a meta-learned self-supervised loss, such that each test-time update improves value estimation. By updating sequentially over a trajectory, VITA encodes history into its parameters, addressing the temporal reasoning limitations. To mitigate shortcut learning, we propose a dissimilarity-based sampling strategy that selects semantically diverse segments of the trajectory during training. In real-world robotic manipulation tasks, VITA generalizes from a single training environment to diverse out-of-distribution tasks, environments, and embodiments, outperforming the state-of-the-art zero-shot method using autoregressive VLMs. Furthermore, we demonstrate that VITA’s zero-shot value estimates can be utilized for reward shaping in offline reinforcement learning, resulting in multi-task policies on the Meta-World benchmark that exceed the performance of those trained with the simulation’s fuzzy-logic dense rewards.
@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated strong generalization across diverse tasks and domains by learning from large-scale, unstructured web data without human supervision [radford2021clip, alayrac2022flamingo]. In contrast, despite significant advances in learning generalist policies for robotics [brohan2023rt2, ghosh2024octo] and 3D virtual environments [reed2022generalist], state-of-the-art methods have yet to achieve comparable success due to their reliance on expert demonstrations [ho2016gail]. World models [ha2018world, hafner2021planet, matsuo2022deep] have emerged as a promising solution, enabling agents to learn in simulated environments, with applications spanning robotics [yang2023unisim, bar2024navigationworldmodels, agarwal2025cosmos], autonomous driving [hu2023gaia1], and video games [bruce2024genie]. Despite the potential of world models to generate realistic visual trajectories in diverse, open-ended environments [hughes2024openended, silver2025era], a critical challenge remains: how can agents effectively learn from videos at scale without relying on human supervision?
@@
A line of research explores policy learning directly from expert visual trajectories by inferring actions from visual transitions [edwards2019imitating, baker2022video, schmidt2024learning, zhang2025latent]. These latent actions can be mapped to executable controls in deployment, although this remains a challenging task in practice. In parallel, other work focuses on learning a universal goal-conditioned value function, using it as a zero-shot reward shaping and supervision signal for reinforcement learning (RL) and imitation learning, respectively [chen2021learning, ma2022vip]. In this framework, goal-conditioned value estimation can be formulated as: predicting how far an agent has progressed toward completing a task, based on visual observations and a natural language task description [lee2021generalizable, ma2023liv, ma2024vlm, dashora2025viva].
@@
Prior work has employed pre-trained contrastive VLMs [radford2021clip] for zero-shot reward shaping, leveraging the similarity between task descriptions and visual observations in a shared multimodal representation space to ground progress in semantic context [ma2023liv, baumli2023vlm, rocamonde2024vlm]. However, these methods fail to capture the temporal context needed to disambiguate visually similar states that occur at different stages of a task (e.g., folding vs. unfolding a shirt). In contrast, autoregressive VLMs [alayrac2022flamingo] incorporate temporal context by conditioning on the entire visual trajectory within the prompt; however, they inherit a bias toward monotonically increasing predictions from the chronologically ordered datasets used in pre-training [ma2024vlm]. Both VLM architectures rely on pre-trained representations [ma2023liv, ma2024vlm, baumli2023vlm, rocamonde2024vlm, du2023success] for zero-shot prediction, which limits both their generalization and their temporal reasoning [patraucean2023perception]. To enable generalizable value function learning, recent work has explored large-scale pretraining [alakuijala2024video, ma2023liv], domain-specific fine-tuning [zhang2025rewind], and demo-based reward specification [sontakke2023roboclip].
@@
In this work, we introduce VITA, a zero-shot value function learning method that enhances both generalization and temporal reasoning through test-time adaptation, outperforming the state-of-the-art zero-shot approach [ma2024vlm] on real-world robotic manipulation tasks. Unlike prior methods that rely on large-scale pretraining or expert demonstrations for generalization, VITA adapts online to both the semantic and temporal context of trajectories. At inference, a lightweight adaptation module is updated in negligible time with a gradient step on a meta-learned self-supervised loss [sun2024rnn], such that each test-time update improves value function estimation [finn2017maml]. By updating sequentially over a test trajectory, the value function estimator encodes trajectory history into its parameters [sun2024rnn], thereby addressing the temporal reasoning limitations of prior work. To mitigate shortcut learning [geirhos2020shortcut], we introduce a dissimilarity-based sampling strategy that encourages reliance on semantic cues, supported by empirical evidence. Our evaluation demonstrates the effectiveness of our method across core capabilities of value function estimation: generalization under distribution shifts, differentiation between expert and non-expert trajectories, and reward shaping for offline RL [levine2020offline].
@@
  • We propose VITA, a test-time adaptation method that enhances the generalization and temporal reasoning of contrastive VLMs for zero-shot value function estimation, without requiring task-specific demonstrations or large-scale pretraining.
  • VITA generalizes from a single training environment to diverse out-of-distribution tasks, environments, and embodiments in robotic manipulation, outperforming the state-of-the-art zero-shot method [ma2024vlm].
  • VITA’s zero-shot value estimates for reward shaping yield offline RL policies on the Meta-World benchmark for multi-task learning (MT10) that surpass those trained with the simulation’s fuzzy-logic dense rewards [yu2020meta].
@@ 2.1 Goal-Conditioned Value Functions
In video trajectories, task progress estimation can be viewed as goal-conditioned value function estimation, where the reward reflects task completion. Therefore, we formulate learning a goal-conditioned value function as predicting the degree of task completion from vision–language representations. Formally, the value function is defined as: $V:\mathcal{O}\times\mathcal{G}\rightarrow[0,1]$ which maps an observation $o_{t}\in\mathcal{O}$ and a goal specification $g\in\mathcal{G}$ to a scalar value indicating the predicted progress toward goal completion. We set $V(o_{t};g)=0$ to correspond to the start of the task and $V(o_{t};g)=1$ to indicate completion. Task progress is commonly aligned with temporal position in expert demonstrations, based on the assumption that such trajectories exhibit monotonically increasing progress toward goal completion [lee2021generalizable, ma2024vlm, dashora2025viva]. Given an expert trajectory $\tau=(o_{1},\ldots,o_{T})\sim\pi_{E}$, temporal progress is defined using normalized timestep indices: $V^{\pi_{E}}(o_{t};g)=\frac{t}{T}$ where $T$ is the trajectory length. Therefore, temporal progress provides supervision for learning a goal-conditioned value function $V$.
@@ 2.2 Test-Time Training for Sequence Modeling
Test-time training (TTT) [sun2020ttt] is a test-time adaptation method that treats inference as a self-supervised learning task [goodfellow2016deep], updating model parameters on each test instance without access to labels. While originally proposed for static image classification [sun2020ttt], recent work extends TTT to sequence modeling [sun2024rnn, wang2023tttvideo], where a parametric adaptation module $f_{\text{adapt}}$ is updated at each timestep using a gradient step on a self-supervised loss. In contrast to standard sequence architectures that encode temporal history in activation states, such as in hidden states [hochreiter1997long, cho2014gru] or in cached key–value pairs [vaswani2017attention], TTT encodes history into the parameters of $f_{\text{adapt}}$ via sequential updates in inference. The parameters of $f_{\text{adapt}}$ thereby serve as an implicit memory, encoding temporal history while preserving temporal order. sun2024rnn further propose to meta-learn the self-supervised task, following the gradient-based meta-learning paradigm [finn2017maml]. In this setting [sun2024rnn], the self-supervised loss is parameterized by learnable linear projections and trained to minimize downstream prediction loss after a test-time update, rather than directly minimizing it. In this work, we apply TTT to goal-conditioned value function learning, where temporal context is captured through sequential updates, and the self-supervised task is meta-learned to improve value estimation rather than being predefined a priori.
@@ 3.1.1 Multimodal Input Representation
We use a frozen contrastive vision-language encoder, CLIP [radford2021clip], to extract representations from visual observations and goal descriptions. Given a visual trajectory $\tau=(o_{1},\ldots,o_{T})$ and a language task description $g$, we concatenate their representations at each timestep $t$ to form a sequence of joint multimodal representations $(z_{1},z_{2},\ldots,z_{T})$, where $z_{t}=[\phi_{v}(o_{t});\phi_{g}(g)]\in\mathbb{R}^{2d}$, and $\phi_{v}$ and $\phi_{g}$ denote the visual and language encoders, respectively. CLIP is pre-trained with a contrastive objective to align paired image-text inputs in a shared representation space [radford2021clip]. As a result, representations of visual observations that are semantically closer to goal completion tend to be closer to the representations of the goal description.
@@ 3.1.3 Zero-shot Value Function Estimator
The regression head $h$ is a two-layer multilayer perceptron (MLP), following value-function estimator architectures used in deep reinforcement learning [espeholt2018impala]. The network is trained using normalized progress labels $y_{t}$ from expert demonstrations as described in Section 2.1, with a supervised prediction loss $\ell_{\text{pred}}$ defined as the mean squared error between predicted values $V(z_{t};g)$ and targets $y_{t}$. At inference, the regression head $h$ remains frozen, while the adaptation module $f_{\text{adapt}}$ is updated online using the self-supervised objective. This enables the value function estimator to generalize in a zero-shot setting without requiring task-specific demonstrations.
@@ 3.2 Training Procedure
We train the value function estimator with gradient-based meta-learning, optimizing a self-supervised loss $\ell_{\text{self}}$ such that test-time adaptation improves the supervised prediction loss $\ell_{\text{pred}}$ [sun2024rnn]. Training is performed on diverse sub-trajectories selected via dissimilarity-based sampling, which mitigates shortcut learning and encourages reliance on semantic and temporal cues.
@@ 3.2.1 Meta-Learning
During training, the adaptation module $f_{\text{adapt}}$ is updated online at each timestep using the self-supervised loss $\ell_{\text{self}}$. Following the gradient-based meta-learning paradigm [finn2017maml], we differentiate through these adaptation updates with respect to the supervised prediction loss $\ell_{\text{pred}}$. This optimizes the initialization $\theta_{0}$ of $f_{\text{adapt}}$, the linear projections $P_{K},P_{V},P_{Q}$, and the regression head $h$. Formally, the total training loss combines the supervised prediction loss $\ell_{\text{pred}}$ with the self-supervised loss $\ell_{\text{self}}$, weighted by a scalar $\lambda$. The learned parameters $\theta_{0}$ serve as the initialization for adaptation at inference time. This approach enables the value function estimator to learn to adapt its internal representations at test time to improve value function estimation.
@@ 3.2.2 Dissimilarity-Based Sampling
Expert visual trajectories often contain consecutive frames that are highly redundant, as noted in prior work on video action recognition [wang2018temporal]. Such redundancy can encourage the value function estimator to exploit shortcut cues [geirhos2020shortcut, lee2021gpil] by overfitting to frequently occurring late-stage visual patterns. To mitigate this, we propose a dissimilarity-based sampling strategy that constructs mini-batches from the most visually diverse sub-trajectories within each trajectory. This increases intra-batch variance and acts as a form of importance sampling that favors semantic diversity.
@@ 4.1 Training Setup
We train VITA on expert visual trajectories paired with natural language task descriptions from the BridgeData V2 dataset [walke2023bridgedata], which spans a wide range of manipulation tasks, environments, and robot embodiments. In particular, we use a curated subset consisting of 2,986 expert demonstrations covering pick-and-place manipulation tasks, but it does not include folding, sweeping, or stacking tasks. All demonstrations are collected using a single robot embodiment (WidowX 250) across 4 configurations of the ToyKitchen environment. An additional 287 expert demonstrations are held out as the in-distribution test set, referred to as tk_pnp. Appendix A includes examples from the training set. We use OpenCLIP ViT-B/32 as the frozen backbone for encoding video frames and task descriptions. The training objective combines $\ell_{\text{pred}}$ with $\ell_{\text{self}}$, weighted by $\lambda_{\text{self}}=0.5$. During training, we use dissimilarity-based sampling with window size $w_{\text{tr}}=8$, number of sub-trajectories $k=8$, and stride $s=1$. At test time, we use one gradient step ($t_{\mathrm{ep}}=1$) on $\lambda_{\text{self}}$ with learning rate $\eta=0.1$. Further details and hyperparameter sweeps are reported in Appendix D.

### v5

# arXiv:2506.10085v5[cs.CV] 27 Feb 2026

@@ Abstract
Vision–Language Models (VLMs) show promise as zero-shot goal-conditioned value functions, but their frozen pre-trained representations limit generalization and temporal reasoning. We introduce VITA, a zero-shot value function learning method that enhances both capabilities via test-time adaptation. At inference, a lightweight adaptation module is updated via a gradient step on a meta-learned self-supervised loss, such that each test-time update improves value estimation. By updating sequentially over a trajectory, VITA encodes history into its parameters, addressing the temporal reasoning limitations. To mitigate shortcut learning, we propose a dissimilarity-based sampling strategy that selects semantically diverse segments of the trajectory during training. In real-world robotic manipulation tasks, VITA generalizes from a single training environment to diverse out-of-distribution tasks, environments, and embodiments, outperforming the state-of-the-art zero-shot method using autoregressive VLMs. Furthermore, we demonstrate that VITA’s zero-shot value estimates can be utilized for reward shaping in offline reinforcement learning, resulting in multi-task policies on the Meta-World benchmark that exceed the performance of those trained with the simulation’s fuzzy-logic dense rewards. Project website: https://chziakas.github.io/vita/
@@ 1 Introduction
Vision-Language Models (VLMs) have demonstrated strong generalization across diverse tasks and domains by learning from large-scale, unstructured web data without human supervision (Radford et al., 2021; Alayrac et al., 2022). In contrast, despite significant advances in learning generalist policies for robotics (Brohan et al., 2023; Kim et al., 2024) and 3D virtual environments (Reed et al., 2022; SIMA Team et al., 2024), state-of-the-art methods have yet to achieve comparable success due to their reliance on expert demonstrations (Ho and Ermon, 2016). World models (Ha and Schmidhuber, 2018; Hafner et al., 2021; Matsuo et al., 2022) have emerged as a promising solution, enabling agents to learn in simulated environments, with applications spanning robotics (Yang et al., 2023; Agarwal et al., 2025), autonomous driving (Hu et al., 2024), and video games (Bruce et al., 2024). Despite the potential of world models to generate realistic visual trajectories in diverse, open-ended environments (Hughes et al., 2024; Silver and Sutton, 2025), a critical challenge remains: how can agents effectively learn from videos at scale without relying on human supervision?
@@
A line of research explores policy learning directly from expert visual trajectories by inferring actions from visual transitions (Edwards et al., 2019; Baker et al., 2022; Schmidt and Jiang, 2024; Zhang et al., 2025a). These latent actions can be mapped to executable controls in deployment, although this remains a challenging task in practice. In parallel, other work focuses on learning a universal goal-conditioned value function, using it as a zero-shot reward shaping and supervision signal for reinforcement learning (RL) and imitation learning, respectively (Chen et al., 2021; Ma et al., 2022). In this framework, goal-conditioned value estimation can be formulated as: predicting how far an agent has progressed toward completing a task, based on visual observations and a natural language task description (Lee et al., 2021a; Ma et al., 2023b; 2024; Dashora et al., 2025).
@@
Prior work has employed pre-trained contrastive VLMs (Radford et al., 2021) for zero-shot reward shaping, leveraging the similarity between task descriptions and visual observations in a shared multimodal representation space to ground progress in semantic context (Ma et al., 2023b; Baumli et al., 2023; Rocamonde et al., 2024). However, these methods fail to capture the temporal context needed to disambiguate visually similar states that occur at different stages of a task (e.g., folding vs. unfolding a shirt). In contrast, autoregressive VLMs (Alayrac et al., 2022) incorporate temporal context by conditioning on the entire visual trajectory within the prompt; however, they inherit a bias toward monotonically increasing predictions from the chronologically ordered datasets used in pre-training (Ma et al., 2024). Both VLM architectures rely on pre-trained representations (Ma et al., 2023b; 2024; Baumli et al., 2023; Rocamonde et al., 2024; Du et al., 2023) for zero-shot prediction, which limits both their generalization and their temporal reasoning (Pătrăucean et al., 2023). To enable generalizable value function learning, recent work has explored large-scale pre-training (Alakuijala et al., 2024; Ma et al., 2023b), domain-specific fine-tuning (Zhang et al., 2025b), and demo-based reward specification (Sontakke et al., 2023).
@@
In this work, we introduce VITA, a zero-shot value function learning method that enhances both generalization and temporal reasoning through test-time adaptation, outperforming the state-of-the-art zero-shot approach (Ma et al., 2024) on real-world robotic manipulation tasks. Unlike prior methods that rely on large-scale pre-training or expert demonstrations for generalization, VITA adapts online to both the semantic and temporal context of trajectories. At inference, a lightweight adaptation module is updated in negligible time with a gradient step on a meta-learned self-supervised loss (Sun et al., 2024), such that each test-time update improves value function estimation (Finn et al., 2017). By updating sequentially over a test trajectory, the value function estimator encodes trajectory history into its parameters (Sun et al., 2024), thereby addressing the temporal reasoning limitations of prior work. To mitigate shortcut learning (Geirhos et al., 2020), we introduce a dissimilarity-based sampling strategy that encourages reliance on semantic cues, supported by empirical evidence. Our evaluation demonstrates the effectiveness of our method across core capabilities of value function estimation: generalization under distribution shifts, differentiation between expert and non-expert trajectories, and reward shaping for offline RL (Levine et al., 2020).
@@
  • We propose VITA, a test-time adaptation method that enhances the generalization and temporal reasoning of contrastive VLMs for zero-shot value function estimation, without requiring task-specific demonstrations or large-scale pre-training.
  • VITA generalizes from a single training environment to diverse out-of-distribution tasks, environments, and embodiments in robotic manipulation, outperforming the state-of-the-art zero-shot method (Ma et al., 2024).
  • VITA’s zero-shot value estimates for reward shaping yield offline RL policies on the Meta-World benchmark for multi-task learning (MT10) that surpass those trained with the simulation’s fuzzy-logic dense rewards (Yu et al., 2020).
@@ 2.1 Goal-Conditioned Value Functions
In video trajectories, task progress estimation can be viewed as goal-conditioned value function estimation, where the reward reflects task completion. Therefore, we formulate learning a goal-conditioned value function as predicting the degree of task completion from vision–language representations. Formally, the value function is defined as: $V:\mathcal{O}\times\mathcal{G}\rightarrow[0,1]$ which maps an observation $o_{t}\in\mathcal{O}$ and a goal specification $g\in\mathcal{G}$ to a scalar value indicating the predicted progress toward goal completion. We set $V(o_{t};g)=0$ to correspond to the start of the task and $V(o_{t};g)=1$ to indicate completion. Task progress is commonly aligned with temporal position in expert demonstrations, based on the assumption that such trajectories exhibit monotonically increasing progress toward goal completion (Lee et al., 2021a; Ma et al., 2024; Dashora et al., 2025). Given an expert trajectory $\tau=(o_{1},\ldots,o_{T})\sim\pi_{E}$, temporal progress is defined using normalized timestep indices: $V^{\pi_{E}}(o_{t};g)=\frac{t}{T}$ where $T$ is the trajectory length. Therefore, temporal progress provides supervision for learning a goal-conditioned value function $V$.
@@ 2.2 Test-Time Training for Sequence Modeling
Test-time training (TTT) (Sun et al., 2020) is a test-time adaptation method that treats inference as a self-supervised learning task (Goodfellow et al., 2016), updating model parameters on each test instance without access to labels. While originally proposed for static image classification (Sun et al., 2020), recent work extends TTT to sequence modeling (Sun et al., 2024; Wang et al., 2023), where a parametric adaptation module $f_{\text{adapt}}$ is updated at each timestep using a gradient step on a self-supervised loss. In contrast to standard sequence architectures that encode temporal history in activation states, such as in hidden states (Hochreiter and Schmidhuber, 1997; Cho et al., 2014) or in cached key–value pairs (Vaswani et al., 2017), TTT encodes history into the parameters of $f_{\text{adapt}}$ via sequential updates in inference. The parameters of $f_{\text{adapt}}$ thereby serve as an implicit memory, encoding temporal history while preserving temporal order. Sun et al. (2024) further propose to meta-learn the self-supervised task, following the gradient-based meta-learning paradigm (Finn et al., 2017). In this setting (Sun et al., 2024), the self-supervised loss is parameterized by learnable linear projections and trained to minimize downstream prediction loss after a test-time update, rather than directly minimizing it. In this work, we apply TTT to goal-conditioned value function learning, where temporal context is captured through sequential updates, and the self-supervised task is meta-learned to improve value estimation rather than being predefined a priori.
@@ 3 VITA: Zero-Shot Value Functions via Test-Time Adaptation
(a) Training: Learning goal-conditioned value function estimator via meta-learning
(b) Inference: Zero-shot value estimate for OOD test trajectory via test-time adaptation
@@ 3.1.1 Multimodal Input Representation
We use a frozen contrastive vision-language encoder, CLIP (Radford et al., 2021), to extract representations from visual observations and goal descriptions. Given a visual trajectory $\tau=(o_{1},\ldots,o_{T})$ and a language task description $g$, we concatenate their representations at each timestep $t$ to form a sequence of joint multimodal representations $(z_{1},z_{2},\ldots,z_{T})$, where $z_{t}=[\phi_{v}(o_{t});\phi_{g}(g)]\in\mathbb{R}^{2d}$, and $\phi_{v}$ and $\phi_{g}$ denote the visual and language encoders, respectively. CLIP is pre-trained with a contrastive objective to align paired image-text inputs in a shared representation space (Radford et al., 2021). As a result, representations of visual observations that are semantically closer to goal completion tend to be closer to the representations of the goal description.
@@ 3.1.3 Zero-shot Value Function Estimator
The regression head $h$ is a two-layer multilayer perceptron (MLP), following value-function estimator architectures used in deep reinforcement learning (Espeholt et al., 2018). The network is trained using normalized progress labels $y_{t}$ from expert demonstrations as described in Section 2.1, with a supervised prediction loss $\ell_{\text{pred}}$ defined as the mean squared error between predicted values $V(z_{t};g)$ and targets $y_{t}$. At inference, the regression head $h$ remains frozen, while the adaptation module $f_{\text{adapt}}$ is updated online using the self-supervised objective. This enables the value function estimator to generalize in a zero-shot setting without requiring task-specific demonstrations.
@@ 3.2 Training Procedure
We train the value function estimator with gradient-based meta-learning, optimizing a self-supervised loss $\ell_{\text{self}}$ such that test-time adaptation improves the supervised prediction loss $\ell_{\text{pred}}$ (Sun et al., 2024). Training is performed on diverse sub-trajectories selected via dissimilarity-based sampling, which mitigates shortcut learning and encourages reliance on semantic and temporal cues.
@@ 3.2.1 Meta-Learning
During training, the adaptation module $f_{\text{adapt}}$ is updated online at each timestep using the self-supervised loss $\ell_{\text{self}}$. Following the gradient-based meta-learning paradigm (Finn et al., 2017), we differentiate through these adaptation updates with respect to the supervised prediction loss $\ell_{\text{pred}}$. This optimizes the initialization $\theta_{0}$ of $f_{\text{adapt}}$, the linear projections $P_{K},P_{V},P_{Q}$, and the regression head $h$. Formally, the total training loss combines the supervised prediction loss $\ell_{\text{pred}}$ with the self-supervised loss $\ell_{\text{self}}$, weighted by a scalar $\lambda$. The learned parameters $\theta_{0}$ serve as the initialization for adaptation at inference time. This approach enables the value function estimator to learn to adapt its internal representations at test time to improve value function estimation.
@@ 3.2.2 Dissimilarity-Based Sampling
Expert visual trajectories often contain consecutive frames that are highly redundant, as noted in prior work on video action recognition (Wang et al., 2018). Such redundancy can encourage the value function estimator to exploit shortcut cues (Geirhos et al., 2020; Lee et al., 2021b) by overfitting to frequently occurring late-stage visual patterns. To mitigate this, we propose a dissimilarity-based sampling strategy that constructs mini-batches from the most visually diverse sub-trajectories within each trajectory. This increases intra-batch variance and acts as a form of importance sampling that favors semantic diversity.
@@ 4.1 Training Setup
We train VITA on expert visual trajectories paired with natural language task descriptions from the BridgeData V2 dataset (Walke et al., 2023), which spans a wide range of manipulation tasks, environments, and robot embodiments. In particular, we use a curated subset consisting of 2,986 expert demonstrations covering pick-and-place manipulation tasks, but it does not include folding, sweeping, or stacking tasks. All demonstrations are collected using a single robot embodiment (WidowX 250) across 4 configurations of the ToyKitchen environment. An additional 287 expert demonstrations are held out as the in-distribution test set, referred to as tk_pnp. Appendix A includes examples from the training set. We use OpenCLIP ViT-B/32 as the frozen backbone for encoding video frames and task descriptions. The training objective combines $\ell_{\text{pred}}$ with $\ell_{\text{self}}$, weighted by $\lambda_{\text{self}}=0.5$. During training, we use dissimilarity-based sampling with window size $w_{\text{tr}}=8$, number of sub-trajectories $k=8$, and stride $s=1$. At test time, we use one gradient step ($t_{\mathrm{ep}}=1$) using $\lambda_{\text{self}}$ with learning rate $\eta=0.1$. Further details and hyperparameter sweeps are reported in Appendix D.